OpenStudy (anonymous):

Can someone teach me how to solve this problem. A gaseous mixture made from 5.0 g of oxygen and 2.5 g of methane is place is a 15 L container at 25°C. What is the partial pressure of oxygen? a. .26 atm b. .52 atm c. 1.0 atm d. 22.4 atm

OpenStudy (abb0t):

You can solve this by using the ideal gas law, \(\sf \color{blue}{PV = nRT}\) to solve for \(\sf \color{violet}{n}\) n = moles. You know that partial pressure of any gas is dependent on number of moles. That means, find number of moles for both oxygen and methane. Then, add them together. Do you follow me here?

OpenStudy (anonymous):

(5.0 g O2) / (31.99886 g O2/mol) = 0.15626 mol O2 P = nRT / V = (0.15626 mol) x (0.08205746 L atm/K mol) x (25 + 273)K / (15 L) = 0.25 atm

OpenStudy (anonymous):

Notice that the methane doesn't enter into this problem at all. That's how partial pressures work.

OpenStudy (anonymous):

Thank you both :) I understand now.